A crate weighing 500.0 N is pulled along a horizontal sidewalk at constant speed by a rope that makes a 27.0

angle with the sidewalk. If a force of 210.0 N is applied to the rope, what is the coefficient of kinetic friction

between the crate and the sidewalk?



1
Expert's answer
2021-01-30T06:26:19-0500
"ma=F\cos{27}-\mu N=0\\ F\cos{27}=\mu (mg-F\sin{27})
\\ 210\cos{27}=\mu (500-210\sin{27})\\\mu=0.46"